OpenStudy (loser66):

parameterize segment. (1-i) to (1+i) . Please help

OpenStudy (loser66):

If I work as usual, I get 1-i + 2it , t from 0 to 1 but my friend makes it easy by 1 + it , t from -1 to 1 I would like to know how he gets it forward, not backward. I meant if we know the parametric equation, then we can check whether it is right or wrong easily. Assume that we don't know it, like I am now, how can we get it from the given information?

OpenStudy (loser66):



OpenStudy (freckles):

is this not right: v=(1,-1)-(1,1)=(0,-2) so answer is (1,1)+t(0,-2) or (1,1-2t) or 1+(1-2t)i

OpenStudy (freckles):

http://mathinsight.org/line_parametrization_examples I used this as example

OpenStudy (loser66):

I don't get what you mean but for the first line segment from 1-i to 1+i, the parameterized equation is not that

OpenStudy (loser66):

if t from 0 to 1, that is my solution. As I said above, my friend get from -1 to 1 with different equation which is faaaaaaaaaaaaar easier than my way on next step. But I don't know how he gets that way (Note: his equation is perfect right)

OpenStudy (freckles):

could you show how he gets backwards?

OpenStudy (loser66):